# On the Compatibility of Constructive Predicative Mathematics with Weyl’s Classical Predicativity

### v3

@@ 3.3 Homotopy Type Theory
Homotopy Type Theory (HoTT) [53] is an extension of MLTT and is an instance of Voevodsky’s Univalent Foundations. It extends MLTT with the Univalence Axiom and Higher Inductive types, examples of which include quotients of homotopy sets and propositional truncation. The Univalence Axiom is formulated as follows:
the $\mbox{the map}\ \mathsf{idtoeqv}:(A=_{\mathcal{U}}B)\ \rightarrow\ (A\simeq B)\ \mbox{is an equivalence}$
where $(A\simeq B)$ is the type of equivalences between $A$ and $B$. The map $\mathsf{idtoeqv}$ and the notion of equivalence $\simeq$ are defined as in Section 2 of [53]. Among the consequences of this axiom, there are certain extensional principles (e.g. function extensionality) that are not available in intensional MLTT.
A significant feature of HoTT is allowing for a representation of logical notions alternative to the propositions-as-types paradigm. The key notion here is that of h-proposition, where a type $P$ is a h-proposition whenever for all $x,y:P$, $x=_{P}y$ holds, and it is analogous to the notion of mono type in the internal theory of topos in [23]. Further, a type $A$ is a h-set if the associated identity type $=_{A}$ is a h-proposition.
The logic of h-propositions behaves differently from the propositions-as-types logic and is similar to the internal logic of elementary topoi or of regular categories as presented in [23]. In particular, the existential quantifier is not identified with the $\Sigma$-type any longer. Rather, it is identified with the propositional truncation of the $\Sigma$:
@@
Observe that the propositional truncation of a type is always a h-proposition. Indeed the constructors for the propositional truncation of a type $A$ are the following: if $a:A$, then $|\ a\ |:||\ A\ ||$ and $x=_{||A||}y$ holds for all $x,y:||\ A\ ||$. Moreover, it holds that if $P$ is a h-proposition, then it is equivalent to its truncation.
@@
with $A,B$ h-sets and $R(x,y)$ h-proposition for $x:A,y:B$, as shown in [53]. Indeed, in HoTT it is only possible to derive the axiom of unique choice AC! formulated as follows
@@
The proof of AC! in [44] is a special case of a result that holds in the internal dependent type theory $\mathcal{T}_{reg}$ of any regular category [23]. Following [23], observe that the propositional truncation of a $\Sigma$-type $\Sigma_{(x:A)}\ P(x)$ can be equivalently expressed in $\mathcal{T}_{reg}$ as a quotient, namely the quotient $\Sigma_{(x:A)}\ P(x)/\mathbf{1}$ over the unit type identifying all the proof-terms. Then, the proof in [44] corresponds to the argument in [23] relying on the fact that unique existence can be written as $\Sigma_{(y:B)}\ R(x,y)$ together with the usual uniqueness condition. In this case, $\Sigma_{(y:B)}\ R(x,y)$ is a mono type, according to the terminology introduced in [23], and is equivalent to its quotient over the unit type, and hence we can use $\Sigma$-projections to prove AC!.
A crucial feature of HoTT is the presence of quotient sets as higher inductive types. This means that Cauchy real numbers à la Bishop can be formalized as a quotient of the type of Cauchy sequences under the equivalence relation $=_{\mathbb{R}_{C}}$ without resorting to setoids. In particular, they form a set in HoTT.
@@
In any case, it must be noticed that real numbers à la Cauchy in HoTT are defined in a different way from the usual type-theoretic definition of Cauchy reals or Bishop reals. Indeed, they are defined as a higher inductive type $\mathbb{R}_{\mathsf{c}}$ ([53], see Sec. 11.3) capturing the idea that Cauchy reals can be regarded as the free complete metric space generated by the rationals and they form an h-set.
@@
For the presentation of ${\bf HA^{\omega}}$, we refer to the standard reference [52].
@@
The language of ${\bf HA^{\omega}}$ includes countably infinite variables $x^{\sigma},y^{\sigma},z^{\sigma},\ldots$ for each type symbol $\sigma\in\mathrm{T}$. A binary predicate for equality $=_{\sigma}$ at each type $\sigma\in\mathrm{T}$ and an application operator $\mathsf{Ap}_{\sigma,\tau}$ for all $\sigma,\tau\in\mathrm{T}$. Furthermore, the language comprehends constants for zero and successor, recursor, pairing and projections and the combinators ${\bf k}$ and ${\bf s}$. Note that the logical implication is denoted with the symbol $\supset$ to avoid confusion with the symbol $\rightarrow$ used to denote function types.
@@
the resulting theory is conservative over HA. This result is known as Goodman’s Theorem [19].
Another significant fact is that ${\bf HA^{\omega}}$ is equiconsistent with its classical version by a double negation translation. This result can be found in [51]. However, this result does not extend to the case of ${\bf HA^{\omega}}+{\bf AC}$. Indeed, it is possible to show that ${\bf HA^{\omega}}+{\bf AC}+{\bf LEM}$, where LEM is the law of excluded middle $\phi\vee\neg\phi$ for all formulas, is strong enough to derive the second-order comprehension axiom and thus to interpret full second-order classical arithmetic ${\bf PA^{2}}$, which is an impredicative theory. The first to observe that ${\bf HA^{\omega}}$ extended with the Axiom of Choice and classical logic allows one to derive the second-order comprehension principle was Spector in [49]. We remark that already the internal number-theoretic rule of unique choice is sufficient to carry out the proof.
@@
We call the rule displayed above internal rule of unique choice to contrast it with the following rule of unique choice: if ${\bf HA^{\omega}}\ \vdash\ \forall x^{\sigma}\ \exists!y^{\tau}\ \phi(x,y)$, then there exists a function $f^{\sigma\ \rightarrow\ \tau}$ such that ${\bf HA^{\omega}}\ \vdash\ \forall x^{\sigma}\ \phi(x,fx)$, where the existence of the choice function is stated externally. Admittedly, in ${\bf HA^{\omega}}$ the two rules are equivalent by the results in [51]. In general, this equivalence holds for all the systems that satisfy the existence property, but it cannot be stated in general. For instance, it does not hold for classical systems.
@@ 5 The incompatibility of MLTT, HoTT and CZF with Weyl’s predicativity
where $A$ is any type, $N_{o}$ is the empty set interpreting the falsum constant, $+$ is the symbol for the disjoint sum type and $\rightarrow$ is the symbol for the function type. Then in MLTT+LEM we can reproduce the argument showing the incompatibility of HAω with Weyl’s predicativity:
@@
It follows from Prop. 5.2 and Cor. 4.4. ∎
@@
As in the case of MLTT, we can reproduce the same proof of Prop.4.3 but with a significant difference. The law of excluded middle cannot be formulated according to the propositions-as-types identifications for all types in HoTT, because this would yield a contradiction in presence of Univalence (see Ch.3 [53]). However, it is still possible to formulate ${\bf LEM}$ for all those types that are h-propositions as
@@
It follows from Prop. 5.5 and Cor. 4.4. ∎
@@
When HoTT is extended with LEM for h-propositions, it is possible to show that the type of h-propositions within any universe $\mathcal{U}$, that is $\mathsf{Prop}_{\mathcal{U}}$, is equivalent (and hence equal by Univalence) to the type of Booleans 2 in $\mathcal{U}$. Therefore, 2 behaves like a subobject classifier and we can prove that the category of h-sets, as defined in [53], is an elementary topos.
@@
It follows from Prop. 5.9 and Cor. 4.4. ∎
It is worth noting that in the case of CZF, the incompatibility with classical predicativity can be shown by a very direct and simple argument proved by Aczel in [1]:
@@ 6 A possible way out: the Minimalist Foundation
MF  is a two-level foundation for constructive mathematics, that was first conceived in [34] and then fully developed in [24]. It consists of an intensional level, called mTT, and an extensional one, called emTT, together with an interpretation of the latter into the first, as defined in [24].
The two-level structure facilitates the compatibility of MF with the most relevant constructive and classical foundations as shown in [24]: the intensional level mTT is compatible with MLTT  and Coquand-Paulin’s Calculus of Inductive Constructions [11], while the extensional level is compatible with CZF, HoTT and the internal theory of elementary topoi in [22, 23]. It is a remarkable property that both levels of MF are compatible with HoTT as shown in [9].
Moreover, both levels of MF extend a version of Martin-Löf’s type theory with a primitive notion of proposition: mTT extends the intensional type theory in [40], while emTT extends the extensional version presented in [39].
@@
The set-constructors of mTT and emTT  include those of first order Martin-Löf’s type theory, respectively as presented in [40] and [39]. Hence we have the following set constructors: the empty set $N_{0}$, the singleton set $N_{1}$, the set of lists $List(A)$ over a set $A$, hence the set of natural numbers $\mathsf{N}$, the indexed sum and the dependent product of the family of sets $B(x)\ set\ [x\in A]$ denoted respectively as $\Sigma_{x\in A}\ B(x)$ and $\Pi_{x\in A}\ B(x)$, the disjoint sum $A+B$ of the set $A$ with the set $B$. In the case of the extensional level emTT  we have that sets are closed under effective quotients $A/R$ over a set $A$, provided that $R$ is a small equivalence relation $R(x,y)\ prop_{s}\ [x\in A,y\in A]$. Such quotient constructors are not included in the intensional level mTT. Furthermore, both mTT and emTT include small propositions $\phi\ prop_{s}$ regarded as sets of their proofs and defined are those propositions closed under intuitionistic connectives, quantifiers, and equalities restricted to sets.
@@
Both propositions of mTT and emTT include standard connectives and quantifiers. However, they differ in the way the rules for propositional equality are given. Indeed, mTT-propositions include an intensional propositional equality defined as in [40], except for the fact that the elimination rule is restricted to propositions. While emTT-propositions include an extensional propositional equality with an elimination rule formulated like that in [39].
@@
It is relevant for our discussion to remark that elimination rules for propositions act only towards propositions and not towards collections. The same applies to small propositions. In this way, mTT and emTT do not generally validate choice principles. In particular, even unique choice cannot be derived in both levels of MF, thanks to results in [30, 31, 25] and the non-derivability of unique choice in Coquand’s Calculus of Constructions CoC, as first proved in [50]. This is in contrast with what happens in MLTT and in HoTT, as we have seen in Section 3. The fact that choice principles are not validated in MF is important in connection with the notion of compatibility introduced in Section 5. In this way, MF turns out to be also compatible with theories where choice does not generally hold, such as CoC.
@@
$q\in\exists_{f\in A\rightarrow B}\ \forall_{x\in A}\ R(\,x,\mathsf{Ap}(f,x)\,)$
@@
Nevertheless, it is worth pointing out that without relying on the existence property, we can show the non-validity of the internal rule of unique choice following the same argument in [25] since this can also be exported to extensions of mTT, not validating the existence property, such as its classical version:
@@
The proof of Prop.3 in [25] can be easily adapted to the case of the internal rule of unique choice. ∎
Now, thanks to the model in [50], we deduce from Prop. 6.3 that
@@
For a proof, we refer to [33]. ∎
Furthermore, from [24, 20] we know a proof-theoretic upper bound for MF:
@@
First, observe that $\bf AC!_{\mathsf{N},\mathsf{N}}$ implies the number-theoretic internal rule of unique choice. The claim then follows from Proposition 6.7 and Cor. 4.4. ∎
@@
It is sufficient to define a translation from ACA to emTT as follows. First-order variables of ACA are translated as variables of type $\mathsf{N}$ in emTT. The constants for zero and the successor are interpreted by the corresponding constructors of type $\mathsf{N}$ in emTT. Second-order variables for sets of natural numbers are interpreted as propositional functions $X\in\mathsf{N}\rightarrow\mathcal{P}(1)$ in emTT and the atomic formulas of the form $t\in X$ are translated as applications of these functions $X(t)$.
In general, first-order formulas of ACA are interpreted as small propositions of emTT with quantifiers and equalities restricted to $\mathsf{N}$. While second-order formulas are interpreted by propositions of emTT quantifying over collections.
Peano axioms can be derived from the rules for natural numbers in emTT and the presence of power-collections, see [24]. The induction scheme of ACA follows from the elimination rule for $\mathsf{N}$ which acts towards all types including all propositions. Finally, the arithmetical comprehension axiom is validated as follows: given an arithmetical formula $\phi(x)$ in ACA, we know that it is interpreted as a small proposition in emTT. Then we can introduce $[\phi](x)\in[x\in\mathsf{N}]$ in $\mathcal{P}(1)$ and we get $X\ {\ :\equiv\ }\lambda x.\phi(x)\in\mathsf{N}\rightarrow\mathcal{P}(1)$. Therefore, it follows that there exists $X$ such that for all $n\in\mathsf{N}$, $X(x)\leftrightarrow\phi$. ∎
Hence, from prop. 6.10 it follows that the classical version of MF can formalize Weyl’s classical predicative reformulation of analysis in Das Kontinuum following Feferman’s analysis in [17, 18].
@@
A direct proof of this was given in [26] but it follows also from the predicativity of the classical version of MF (see also [33]). If real numbers form a set in MF then this is so in the classical version of MF and every interval of real numbers form a set too. Since classically, the power-collection of natural numbers can be put in bijection with an interval of real numbers, then it would follow that impredicative definitions can be employed in the classical version of MF contrary to its predicativity established thanks in prop.6.7. ∎
Despite the absence of choice principles and quantification over all real numbers to form new sets, a development of mathematics, and in particular of topology and analysis, within emTT is possible if point-free methods are adopted as advocated by Martin-Löf and Sambin with the introduction of Formal Topology in [45]. Indeed, in emTT the following point-free topologies are definable:
@@
The formalization of the point-free topology of the Cantor space can be carried on in emTT as in section 2.1 of [54]. Then, observe that the point-free topology of real numbers can be defined as in Def.3 of [7] in terms of another cover generated by a finite set of axioms that can be represented in emTT with a completely analogous proof to that just mentioned for the point-free topology of Cantor space. ∎
Therefore, one can carry on in MF the development of analysis as pursued in [7, 42, 43] (see also [21]) and that of Positive Topology in [47] and loc.cit..

### v4

@@ 3.3 Homotopy Type Theory
Homotopy Type Theory (HoTT) [hottbook] is an extension of MLTT and is an instance of Voevodsky’s Univalent Foundations. It extends MLTT with the Univalence Axiom and Higher Inductive types, examples of which include quotients of homotopy sets and propositional truncation. The Univalence Axiom is formulated as follows:
$\mbox{the map}\ \mathsf{idtoeqv}:(A=_{\mathcal{U}}B)\ \rightarrow\ (A\simeq B)\ \mbox{is an equivalence}$
where $(A\simeq B)$ is the type of equivalences between $A$ and $B$. The map $\mathsf{idtoeqv}$ and the notion of equivalence $\simeq$ are defined as in Section 2 of [hottbook]. Among the consequences of this axiom, there are certain extensional principles (e.g. function extensionality) that are not available in intensional MLTT.
A significant feature of HoTT is allowing for a representation of logical notions alternative to the propositions-as-types paradigm. The key notion here is that of h-proposition, where a type $P$ is a h-proposition whenever for all $x,y:P$, $x=_{P}y$ holds, and it is analogous to the notion of mono type in the internal theory of topos in [Maietti05]. Further, a type $A$ is a h-set if the associated identity type $_{A}$ is a h-proposition.
The logic of h-propositions behaves differently from the propositions-as-types logic and is similar to the internal logic of elementary topoi or of regular categories as presented in [Maietti05]. In particular, the existential quantifier is not identified with the $\Sigma$-type any longer. Rather, it is identified with the propositional truncation of the $\Sigma$:
@@
[…]
@@
with $A,B$ h-sets and $R(x,y)$ h-proposition for $x:A,y:B$, as shown in [hottbook]. Indeed, in HoTT it is only possible to derive the axiom of unique choice AC! formulated as follows
@@
The proof of AC! in [rijke2015sets] is a special case of a result that holds in the internal dependent type theory $\mathcal{T}_{reg}$ of any regular category [Maietti05]. Following [Maietti05], observe that the propositional truncation of a $\Sigma$-type $\Sigma_{(x:A)}\ P(x)$ can be equivalently expressed in $\mathcal{T}_{reg}$ as a quotient, namely the quotient $\Sigma_{(x:A)}\ P(x)/\mathbf{1}$ over the unit type identifying all the proof-terms. Then, the proof in [rijke2015sets] corresponds to the argument in [Maietti05] relying on the fact that unique existence can be written as $\Sigma_{(y:B)}\ R(x,y)$ together with the usual uniqueness condition. In this case, $\Sigma_{(y:B)}\ R(x,y)$ is a mono type, according to the terminology introduced in [Maietti05], and is equivalent to its quotient over the unit type, and hence we can use $\Sigma$-projections to prove AC!.
A crucial feature of HoTT is the presence of quotient sets as higher inductive types. This means that Cauchy real numbers à la Bishop can be formalized as a quotient of the type of Cauchy sequences under the equivalence relation $_{\mathbb{R}_{C}}$ without resorting to setoids. In particular, they form a set in HoTT.
@@
In any case, it must be noticed that real numbers à la Cauchy in HoTT are defined in a different way from the usual type-theoretic definition of Cauchy reals or Bishop reals. Indeed, they are defined as a higher inductive type $\mathbb{R}_{\mathsf{c}}$ ([hottbook], see Sec. 11.3) capturing the idea that Cauchy reals can be regarded as the free complete metric space generated by the rationals and they form an h-set.
@@
For the presentation of $\bf HA^{\omega}$}, we refer to the standard reference [TVD88].
@@
The language of $\bf HA^{\omega}$} includes countably infinite variables $x^{\sigma},y^{\sigma},z^{\sigma},\ldots$ for each type symbol $\sigma\in\mathrm{T}$. A binary predicate for equality $_{\sigma}$ at each type $\sigma\in\mathrm{T}$ and an application operator $\mathsf{Ap}_{\sigma,\tau}$ for all $\sigma,\tau\in\mathrm{T}$. Furthermore, the language comprehends constants for zero and successor, recursor, pairing and projections and the combinators $\bf k$} and $\bf s$}. Note that the logical implication is denoted with the symbol $\supset$ to avoid confusion with the symbol $\rightarrow$ used to denote function types.
@@
the resulting theory is conservative over HA. This result is known as Goodman’s Theorem [goodman1978].
Another significant fact is that $\bf HA^{\omega}$} is equiconsistent with its classical version by a double negation translation. This result can be found in [troelstra1973]. However, this result does not extend to the case of ${\bf HA^{\omega}}+{\bf AC}$. Indeed, it is possible to show that ${\bf HA^{\omega}}+{\bf AC}+{\bf LEM}$, where LEM is the law of excluded middle $\phi\vee\neg\phi$ for all formulas, is strong enough to derive the second-order comprehension axiom and thus to interpret full second-order classical arithmetic $\bf PA^{2}$}, which is an impredicative theory. The first to observe that $\bf HA^{\omega}$} extended with the Axiom of Choice and classical logic allows one to derive the second-order comprehension principle was Spector in [spector1962provably]. We remark that already the internal number-theoretic rule of unique choice is sufficient to carry out the proof.
@@
We call the rule displayed above internal rule of unique choice to contrast it with the following rule of unique choice: if ${\bf HA^{\omega}}\ \vdash\ \forall x^{\sigma}\ \exists!y^{\tau}\ \phi(x,y)$, then there exists a function $f^{\sigma\ \rightarrow\ \tau}$ such that ${\bf HA^{\omega}}\ \vdash\ \forall x^{\sigma}\ \phi(x,fx)$, where the existence of the choice function is stated externally. Admittedly, in $\bf HA^{\omega}$} the two rules are equivalent by the results in [troelstra1973]. In general, this equivalence holds for all the systems that satisfy the existence property, but it cannot be stated in general. For instance, it does not hold for classical systems.
@@ 5 The incompatibility of MLTT, HoTT and CZF with Weyl’s predicativity
where $A$ is any type, $N_{o}$ is the empty set interpreting the falsum constant, ++ is the symbol for the disjoint sum type and $\rightarrow$ is the symbol for the function type. Then in MLTT+LEM we can reproduce the argument showing the incompatibility of HAω with Weyl’s predicativity:
@@
It follows from Prop. 5.2 and Cor. $4.$ ∎
@@
As in the case of MLTT, we can reproduce the same proof of Prop.4.3 but with a significant difference. The law of excluded middle cannot be formulated according to the propositions-as-types identifications for all types in HoTT, because this would yield a contradiction in presence of Univalence (see Ch.3 [hottbook]). However, it is still possible to formulate $\bf LEM$} for all those types that are h-propositions as
@@
It follows from Prop. 5.5 and Cor. $4.$ ∎
@@
When HoTT is extended with LEM for h-propositions, it is possible to show that the type of h-propositions within any universe $\mathcal{U}$, that is $\mathsf{Prop}_{\mathcal{U}}$, is equivalent (and hence equal by Univalence) to the type of Booleans 2 in $\mathcal{U}$. Therefore, 2 behaves like a subobject classifier and we can prove that the category of h-sets, as defined in [hottbook], is an elementary topos.
@@
It follows from Prop. 5.9 and Cor. $4.$ ∎
It is worth noting that in the case of CZF, the incompatibility with classical predicativity can be shown by a very direct and simple argument proved by Aczel in [aczel78]:
@@ 6 A possible way out: the Minimalist Foundation
MF is a two-level foundation for constructive mathematics, that was first conceived in [ms05] and then fully developed in [m09]. It consists of an intensional level, called mTT, and an extensional one, called emTT, together with an interpretation of the latter into the first, as defined in [m09].
The two-level structure facilitates the compatibility of MF with the most relevant constructive and classical foundations as shown in [m09]: the intensional level mTT is compatible with MLTT and Coquand-Paulin’s Calculus of Inductive Constructions [Coqpau], while the extensional level is compatible with CZF, HoTT and the internal theory of elementary topoi in [LS1986, Maietti05]. It is a remarkable property that both levels of MF are compatible with HoTT as shown in [cm2024].
Moreover, both levels of MF extend a version of Martin-Löf’s type theory with a primitive notion of proposition: mTT extends the intensional type theory in [nordstrom1990programming], while emTT extends the extensional version presented in [martin-lof:bibliopolis].
@@
The set-constructors of mTT and emTT include those of first order Martin-Löf’s type theory, respectively as presented in [nordstrom1990programming] and [martin-lof:bibliopolis]. Hence we have the following set constructors: the empty set $N_{0}$, the singleton set $N_{1}$, the set of lists $List(A)$ over a set $A$, hence the set of natural numbers $\mathsf{N}$, the indexed sum and the dependent product of the family of sets $B(x)\ set\ [x\in A]$ denoted respectively as $\Sigma_{x\in A}\ B(x)$ and $\Pi_{x\in A}\ B(x)$, the disjoint sum $A+B$ of the set $A$ with the set $B$. In the case of the extensional level emTT  we have that sets are closed under effective quotients $A/R$ over a set $A$, provided that $R$ is a small equivalence relation $R(x,y)\ prop_{s}\ [x\in A,y\in A]$. Such quotient constructors are not included in the intensional level mTT. Furthermore, both mTT and emTT include small propositions $\phi\ prop_{s}$ regarded as sets of their proofs and defined are those propositions closed under intuitionistic connectives, quantifiers, and equalities restricted to sets.
@@
Both propositions of mTT and emTT include standard connectives and quantifiers. However, they differ in the way the rules for propositional equality are given. Indeed, mTT-propositions include an intensional propositional equality defined as in [nordstrom1990programming], except for the fact that the elimination rule is restricted to propositions. While emTT-propositions include an extensional propositional equality with an elimination rule formulated like that in [martin-lof:bibliopolis].
@@
It is relevant for our discussion to remark that elimination rules for propositions act only towards propositions and not towards collections. The same applies to small propositions. In this way, mTT and emTT do not generally validate choice principles. In particular, even unique choice cannot be derived in both levels of MF, thanks to results in [qu12, MR16, Ma17] and the non-derivability of unique choice in Coquand’s Calculus of Constructions CoC, as first proved in [st92]. This is in contrast with what happens in MLTT and in HoTT, as we have seen in Section 3. The fact that choice principles are not validated in MF is important in connection with the notion of compatibility introduced in Section 5. In this way, MF turns out to be also compatible with theories where choice does not generally hold, such as CoC.
@@
$q\in\exists_{f\in A\rightarrow B}\ \forall_{x\in A}\ R(\,x,\mathsf{Ap}(f,x)\,)\ \$
@@
Nevertheless, it is worth pointing out that without relying on the existence property, we can show the non-validity of the internal rule of unique choice following the same argument in [Ma17] since this can also be exported to extensions of mTT, not validating the existence property, such as its classical version:
@@
The proof of Prop.3 in [Ma17] can be easily adapted to the case of the internal rule of unique choice. ∎
Now, thanks to the model in [st92], we deduce from Prop. 6.3 that
@@
For a proof, we refer to [MSdoubleneg]. ∎
Furthermore, from [m09, IMMS] we know a proof-theoretic upper bound for MF:
@@
First, observe that $\bf AC!_{\mathsf{N},\mathsf{N}}$ implies the number-theoretic internal rule of unique choice. The claim then follows from Proposition 6.7 and Cor. $4.$ ∎
@@
It is sufficient to define a translation from ACA to emTT as follows. Numerical variables of ACA are translated as variables of type $\mathsf{N}$ in emTT. Set variables of ACA are translated as terms of type $\mathcal{P}(\mathsf{N})$. The symbols for zero and the successor in ACA are interpreted by the corresponding constructors for $\mathsf{N}$ in emTT. Atomic formulas of the form $s=t$ are translated as propositional equalities over $\mathsf{N}$ and those of the form $t\in X$ as small propositions $t\ \varepsilon\ X$ [^3] This notation can be defined in emTT. Given $X\in\mathcal{P}(A)$ and $t\in A$, we define $t\ \varepsilon\ X$ as $\mathsf{Eq}(\mathcal{P}(1),X(t),[\top])$. Moreover, it holds that $t\ \varepsilon\ \{x\in A\mid\phi(x)\}\leftrightarrow\phi(t)$ is true. . In general, first-order formulas of ACA are translated as small propositions with quantifiers ranging over $\mathsf{N}$, while second-order formulas are translated as propositions with quantifiers ranging over the collection $\mathcal{P}(\mathsf{N})$.
Peano axioms can be derived in emTT thanks to the rules for $\mathsf{N}$ and the presence of $\mathcal{P}(1)$. In particular, the induction scheme of ACA follows from the elimination rule for $\mathsf{N}$, which acts towards all collections, including propositions.
Finally, the arithmetical comprehension axiom is validated as follows: given a small proposition $\phi(x)\ [x\in\mathsf{N}]$, we can introduce $\phi(x)]\in\mathcal{P}(1)\ [x\in\mathsf{N}]$ and then form $\{x\in\mathsf{N}\ \mid\ \phi(x)\}\in\mathcal{P}(\mathsf{N})$. Then, it follows that there is an $X\in\mathcal{P}(\mathsf{N})$ such that for all $x\in\mathsf{N}$, $x\ \varepsilon\ X\ \leftrightarrow\ \phi(x)$ holds. ∎
Hence, from prop. 6.10 it follows that the classical version of MF can formalize Weyl’s classical predicative reformulation of analysis in Das Kontinuum following Feferman’s analysis in [feferman1998light, feferman2000significance].
@@
A direct proof of this was given in [ober2020] but it follows also from the predicativity of the classical version of MF (see also [MSdoubleneg]). If real numbers form a set in MF then this is so in the classical version of MF and every interval of real numbers form a set too. Since classically, the power-collection of natural numbers can be put in bijection with an interval of real numbers, then it would follow that impredicative definitions can be employed in the classical version of MF contrary to its predicativity established thanks in prop.6.7. ∎
Despite the absence of choice principles and quantification over all real numbers to form new sets, a development of mathematics, and in particular of topology and analysis, within emTT is possible if point-free methods are adopted as advocated by Martin-Löf and Sambin with the introduction of Formal Topology in [sambin1987]. Indeed, in emTT the following point-free topologies are definable:
@@
The formalization of the point-free topology of the Cantor space can be carried on in emTT as in section 2.1 of [Valentini-bar]. Then, observe that the point-free topology of real numbers can be defined as in Def.3 of [CN95] in terms of another cover generated by a finite set of axioms that can be represented in emTT with a completely analogous proof to that just mentioned for the point-free topology of Cantor space. ∎
Therefore, one can carry on in MF the development of analysis as pursued in [CN95, Palmgren05, pal07] (see also [kawai2023]) and that of Positive Topology in [sambinbook] and loc.cit..